The following are the dimensions of a metallic plate ; Length = 15.6= 0.1cm, Width = 5.51= 0.02cm, thickness =0.84= 0.01cm and mass = 12.512= 0.001g. Calculate the density of the plate and indicate the order accuracy.
1
Expert's answer
2021-10-24T18:25:30-0400

"L=0.1 \;cm \\
W = 0.02 \;cm \\
T = 0.01 \;cm \\
m= 0.001 \;g \\
V = LWT \\
V = 0.1 \times 0.02 \times 0.01 = 2.0 \times 10^{-4} \;cm^3 \\
Density = ρ = \frac{m}{V} = \frac{0.001}{2.0 \times 10^{-4}} = 5.0 \;g/cm^3"